OpenStudy (anonymous):

Write as a product of linear factors: f(x)= x^4 -100 ??

OpenStudy (anonymous):

start by factoring the difference of two squares \[x^4-100=(x^2+10)(x^2-10)\]

OpenStudy (anonymous):

then for the second part repeat to get \[(x^2+10)(x+\sqrt{10})(x-\sqrt{10})\]

OpenStudy (anonymous):

if you need linear factors you need complex numbers because the first part does not factor over the integers. you would get \[(x+\sqrt{10}i)(x-\sqrt{10}i)(x+\sqrt{10})(x-\sqrt{10})\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw